# Dissipation-Induced Steady States in Topological Superconductors: Mechanisms and Design Principles

### v1

@@ Abstract
The search for conditions supporting degenerate steady states in nonequilibrium topological superconductors is important for advancing dissipative quantum engineering, a field that has attracted significant research attention over the past decade. In this study, we address this problem by investigating topological superconductors hosting unpaired Majorana modes under the influence of environmental dissipative fields. Within the Gorini-Kossakowski-Sudarshan-Lindblad framework and the third quantization formalism, we establish a correspondence between equilibrium Majorana zero modes and non-equilibrium kinetic zero modes. We further derive a simple algebraic relation between the numbers of these excitations expressed in terms of hybridization between the single-particle wavefunctions and linear dissipative fields. Based on these findings, we propose a practical recipes how to stabilize degenerate steady states in topological superconductors through controlled dissipation engineering. To demonstrate their applicability, we implement our general framework in the BDI-class Kitaev chain with long-range hopping and pairing terms — a system known to host a robust edge-localized Majorana modes.
@@ 1 Introduction
An important issue of dissipative quantum engineering is the NESS dimension analysis. In the case of the GKSL equation, a series of theorems were proved defining algebraic conditions for the Hamiltonian of the system and jump operators for which the stationary solution is unique [41, 42, 43, 44, 45, 46]. Further interest in the formation of dark spaces led to the emergence of theorems that set sufficient conditions for existence of degenerate NESS. So, for example, it was shown that if the GKSL equation has the so-called strong symmetries, given the existence of unitary operators with which both the Hamiltonian of the system and the jump operators commute, the number of NESS is determined by the number of invariant subspaces of operators of such symmetries [47, 48, 49, 50, 51]. An analog of the Lieb-Schultz-Mattis theorem has recently been proved for open systems with translational invariance and strong $U(1)$ symmetry, which states that it is impossible to implement a single NESS with an odd total charge of the system [52]. We emphasize that the described necessary conditions for the existence of degenerate NESS require a special class of dissipative fields: either reflecting the symmetry of the Hamiltonian, or corresponding to a homogeneous system preserving the total charge.
@@
The outline of the paper is as follows. We formulate our dissipative model in Sec. 2. In Sec. 3 we review the third quantization method in application to the considered model. The conditions for existence of ZKM as well as corresponding strong symmetry are derived in Sec. 4. Our general results applied to Kitaev chain are described in Sec. 5. In Sec. 6 we end the paper with conclusions. Some technical details are delegated to the Appendix 7.
@@ 2.1 The model of dissipative dynamics
In this section, we formulate a model of superconductors that are affected by dissipative fields that do not preserve the number of fermions. As we demonstrate in Appendix 7, coupling to reservoirs can be also described as the action of the dissipative fields. Therefore, in what follows, we do not consider coupling to the reservoirs separately. We assume that the evolution of the density matrix of the system affected by $N_{B}$ dissipative channels satisfies the following GKSL equation:
@@
$\hat{w}_{2l-1}=\hat{c}_{l}+\hat{c}^{+}_{l},\qquad\hat{w}_{2l}=i(\hat{c}_{l}-\hat{c}_{l}^{+}),\qquad l=1,\dots,N.$ (3)
@@ 2.2 Solution for the unitary evolution
$\mathcal{W}=[\,\underline{\chi}_{1},\leavevmode\nobreak\ \underline{\chi}_{2},\ldots,\leavevmode\nobreak\ \underline{\chi}_{2a-1},\leavevmode\nobreak\ \underline{\chi}_{2a},\ldots,\leavevmode\nobreak\ \underline{\chi}_{2N-1},\leavevmode\nobreak\ \underline{\chi}_{2N}\,]\,\in\mathbb{R}^{2N\times 2N}.$ (7)
@@
$\varepsilon_{a}=\underline{\chi}_{2a}^{T}A\,\underline{\chi}_{2a-1},\qquad a=1,\dots,N$ (8)
@@ 3.1 The Fock-Liouville space
$|\,\hat{P}_{\underline{\alpha}}\,\rangle=\hat{P}_{\alpha_{1},\alpha_{2},\ldots,\alpha_{2N}}=2^{-N}\,\hat{w}_{1}^{\alpha_{1}}\,\hat{w}_{2}^{\alpha_{2}}\dots\hat{w}_{2N}^{\alpha_{2N}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{j}\in\{0,\,1\},$ (13)
@@
$\langle\,\hat{P}_{\underline{\beta}}\,|\,\hat{P}_{\underline{\alpha}}\,\rangle=\operatorname{Tr}\,\left(\,\hat{P}^{+}_{\underline{\beta}}\cdot\hat{P}_{\underline{\alpha}}\,\right),$ (14)
@@ 3.2 Mapping of GKSL equation to the Fock-Liouville space
$\mathcal{U}=\frac{1}{\sqrt{2}}\begin{pmatrix}I_{N}&I_{N}\\ -iI_{N}&iI_{N}\\ \end{pmatrix}$. (23)
@@ 3.3 Spectrum of the Liouvillian
$XF+FX^{T}-Y=0\,,\leavevmode\nobreak\ \leavevmode\nobreak\ Y=-Y^{*}=-Y^{T}.$ (25)
The above equation has the only solution for the matrix $F$ provided $\operatorname{Re}\beta_{l}>0$ for all $l$. Physical meaning of $F$ is the non-equilibrium correlation function in Majorana representation: $F_{ij}=\frac{1}{2}\langle\,\left[\,\hat{w}_{i}\,,\,\hat{w}_{j}\,\right]\,\rangle$. Special care is needed if there exists the zero eigen values of $X$, $\beta_{l}=0$, or pairs of purely imaginary eigen values, $\beta_{l}+\beta_{\bar{l}}=0$. In this case the kernel of the operator $\mathcal{R}(F)=XF+FX^{T}$ is non-trivial. As it follows from the Fredholm alternative [89] a solution for $F$ exists if and only if the matrix $Y$ is orthogonal to the kernel of the operator $\bar{\mathcal{R}}(Z)=X^{T}Z+ZX$, i.e. $\operatorname{Tr}\left(Y^{T}Z\right)=0$ for all solutions of the equation $\bar{\mathcal{R}}(Z)=0$, see Theorem 1.1 of Ref. [90]. Let us consider a pair of eigenvalues with zero real part, $\beta_{l}$ and $\operatorname{Im}\beta_{\bar{l}}=-\operatorname{Im}\beta_{l}$. Then we can construct the solution for the equation $\bar{\mathcal{R}}(Z)=0$ as $Z_{km}=U_{kl}(U^{-1}_{lm})^{*}$, where there is no summation over the index $l$. Then we find the following condition for the existence of solutions for $F$: $U^{-1*}YU]_{ll}=0$. Also, the obvious condition should hold $U^{-1*}YU^{-1\,T}]_{ll}=0$. In practice one can use the regularization $X\to X_{\epsilon}=X+\epsilon UU^{-1}$ with $\epsilon\to 0^{+}$ in order to find the solution of Eq. (25). We emphasize that such solution does not contain the information about ZKM and eigen states of $X$ with $\beta_{\bar{l}}=-\beta_{l}$. Also we note that this solution is not unique, since one can add to it the solution $\delta F$ of the homogeneous equation $X\delta F+\delta FX^{T}=0$. We note that one can use any pair of eigen values with $\beta_{\bar{l}}=-\beta_{l}$ to construct the solution of the homogeneous equation as $\delta F^{(l\bar{l})}\sim\underline{U}_{l}\underline{U}_{\bar{l}}^{T}-\underline{U}_{\bar{l}}\underline{U}_{l}^{T}$.
@@
$\!\!\!\!\!\!{V}\,{\mathcal{L}}\,{V}^{-1}=\begin{pmatrix}B&0\\ 0&-B\end{pmatrix}\,,\quad{V}=\begin{pmatrix}U^{-1}&U^{-1}F\\ 0&U^{T}\\ \end{pmatrix}\,,\leavevmode\nobreak\ {V}^{-1}=\begin{pmatrix}U&-F\,(\,U^{-1}\,)^{T}\\ \!0&\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (\,U^{-1}\,)^{T}\\ \end{pmatrix}$. (26)
@@
$\check{\mathcal{L}}=-{4}\sum_{l=1}^{2N}\,\beta_{l}\,\check{b}_{l}^{\times}\check{b}_{l}\,;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \begin{pmatrix}\underline{\check{b}}\\ \,\underline{\check{{b}}}^{\times}\\ \end{pmatrix}={V}\begin{pmatrix}\underline{\check{c}}\\ \underline{\check{{c}}}^{+}\end{pmatrix}=\begin{pmatrix}U^{-1}\,\underline{\check{c}}\,+\,U^{-1}\,F\,\underline{\check{c}}^{+}\\ \,U^{T}\,\underline{\check{{c}}}^{+}\end{pmatrix}\,$. (27)
Here the third quantized operators $\check{b}$ and $\check{b}^{\times}$ satisfy fermionic commutation relations: $\{\,\check{b}_{l},\,\check{b}_{m}\,\}=\{\,\check{b}^{\times}_{l},\,\check{b}^{\times}_{m}\,\}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \{\,\check{b}_{l},\,\check{b}^{\times}_{m}\,\}=\delta_{lm};\leavevmode\nobreak\ \leavevmode\nobreak\ l,m=1,\ldots,2N.$ We note that $\check{b}^{\times}_{l}\neq\check{b}_{l}^{+}$.
@@
The unit coefficient in front of $|\,\hat{\rho}_{NESS}\,\rangle$ is fixed by the requirement of trace invariance, $\operatorname{Tr}\,\hat{\rho}_{NESS}=1$, while the relations $\alpha_{r}=\alpha_{r}^{*}$ and $a_{p}=a_{p}^{*}$ are related with the requirement of the density matrix to be Hermitian, $\hat{\rho}_{NESS}=\hat{\rho}_{NESS}^{+}$. Therefore, the dark space is parametrized by the set of real numbers $\alpha_{r}$, $a_{p}$, $\ldots$ Dots in Eq. (29) denotes the terms which are constructed in a similar way: two pairs of modes with $\beta_{l}+\beta_{l^{\prime}}=0$ are excited, three pairs and so on.
@@
Such form of the density matrix guarantees the validity of the Wick theorem for the correlation functions in Majorana representation. The hermitian operator $\hat{\mathcal{H}}=(\,i/2\,)\,\underline{\hat{w}}^{T}G\underline{\hat{w}}$ involved in Eq. (30) is interpreted as the effective Hamiltonian of the dissipative system in the non-equilibrium stationary state. We note that the matrix $F$ can be explicitly rewritten as the correlation matrix $F_{ij}=\operatorname{Tr}(\left[\,\hat{w}_{i}\,,\,\hat{w}_{j}\,\right]\hat{\rho}_{G})/2$. We note that information about ZKM do not appear in the spectrum of the matrix $G$ since it is absent in the matrix $F$.
@@ 3.4 Types and robustness of zero modes in closed and open systems
$\delta\langle\,\hat{O}\,\rangle_{\rm iso}=\langle 1|\hat{O}|1\rangle-\langle 0|\hat{\mathcal{O}}|0\rangle\equiv\operatorname{Tr}\left(\,\hat{O}\,\delta\hat{\rho}^{(2)}\,\right)=4\,\underline{\chi}_{1}^{T}O\underline{\chi}_{2}.$ (31)
@@
$\delta\langle\,\hat{O}\,\rangle^{(1)}_{\rm dis}=\operatorname{Tr}\left(\,\hat{O}\,\hat{\rho}^{(1)}\,\right)\equiv 0$ (35)
@@
$\delta\langle\,\hat{O}\,\rangle^{(2)}_{\rm dis}=\operatorname{Tr}\left(\,\hat{O}\,\hat{\rho}^{(2)}\,\right)=\frac{1}{4}\sum_{(p\,\bar{p}):\,\beta_{p}+\beta_{\bar{p}}=0}a_{p}\,\operatorname{Im}\left(\,\underline{U}_{p}^{T}\,O\,\underline{U}_{\bar{p}}\,\right).$ (36)
@@
Thus, we argued that occupation of a single ZKM with $\beta_{r}=0$ leads to an indistinguishable change in the structure of a degenerate non-equilibrium stationary state from the point of view of single-particle physical observables. This property of ZKM is different from the property of zero energy states in the isolated case described in Sec. 2. For the latter it holds only for local operators. Such a difference is probably explained by the fact that the NESS forms as a result of hybridization with the environment. In systems with the multiple ZKM, the hybridization behavior becomes more subtle (see the Sec. 4.3). Nevertheless, the NESS states remain indistinguishable to at least local observables. Therefore, a knowledge of the number of ZKM is crucial for fixing the dimensionality of the dark space in the presence of dissipation.
@@ 3.5 Limit of vanishing dissipation
where we used Eqs. (10) and (24), and the relations $\hat{\rho}^{(1)}\to 0$, $\leavevmode\nobreak\ \underline{\psi}_{2l-1(2l)}\to\underline{\chi}_{2l-1(2l)}$, $\sum_{l}\,U_{l\bar{p}}\,\hat{w}_{l}/2\to\hat{\alpha}_{p}^{+}$ in the limit $\gamma{\to}0^{+}$. Enforcing consistency between Eqs. (39)-(41) yields  [^4] Note that in the limit $\gamma\to 0^{+}$, the matrices $F$ and $X$ have a common eigenbasis defined by the matrix $U$. This corresponds to the situation where the so-called spectral and purity gaps coincide [69]. For finite $\gamma$, however, this property generally no longer holds, and the two gaps characterize distinct features of the system.:
@@ 4 The condition for the existence of zero kinetic modes
4.1 The number of ZKM: algebraic approach
where the matrices $\mathcal{Z}_{0,1,2}$ read
we find for $\beta\to 0$:
$\mathcal{P}(\beta\to 0)\sim\beta^{2N_{M}-\operatorname{rk}\mathcal{Z}_{0}}\,,\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname{rk}\mathcal{Z}_{0}\leq 2\min(N_{B},\,N_{M})\,.$ (51)
where we introduce
Here $\tilde{\mathcal{W}}=[\underline{\chi}_{1},\underline{\chi}_{2},\dots,\underline{\chi}_{2N_{M}-1},\underline{\chi}_{2N_{M}}]\in\mathbb{R}^{2N\times 2N_{M}}$ is nothing but the reduced $\mathcal{W}$, cf. Eq. (7), composed out of $2N_{M}$ first columns, corresponding to the zero modes. We note that the matrix $\mathcal{B}$ has the form of cross Gram matrix built out of two sets of vectors $\underline{\chi}_{a}$, $a=1,\dots,2N_{M}$, and $l_{v}^{r,i}$, $v=1,\dots,N_{B}$. The matrix $\mathcal{B}$ has a transparent meaning. The vectors $\chi_{a}$, $a=1,\dots,2N$, form a basis in $\mathbb{R}^{2N}$. The eigen vectors $\chi_{a}$, $a=1,\dots,2N_{M}$, correspond to the zero mode subspace in $\mathbb{R}^{2N}$. The elements of the matrix $\mathcal{B}$ encodes the part of the dissipative field belonging to the zero mode subspace: $\ell=\mathcal{B}^{T}\tilde{\mathcal{W}}$$\ell_{\perp}$.

### v2

@@ Abstract
The search for conditions supporting degenerate steady states in nonequilibrium topological superconductors is important for advancing dissipative quantum engineering, a field that has attracted significant research attention over the past decade. In this study, we address this problem by investigating topological superconductors hosting unpaired Majorana modes under the influence of environmental dissipative fields. Within the Gorini-Kossakowski-Sudarshan-Lindblad framework and the third quantization formalism, we establish a correspondence between equilibrium Majorana zero modes and non-equilibrium kinetic zero modes. We further derive a simple algebraic relation between the numbers of these excitations expressed in terms of hybridization between the single-particle wavefunctions and linear dissipative fields. Based on these findings, we propose a practical recipes how to stabilize degenerate steady states in topological superconductors through controlled dissipation engineering. To demonstrate their applicability, we implement our general framework in the BDI-class Kitaev chain with long-range hopping and pairing terms – a system known to host a robust edge-localized Majorana modes.
@@ 1 Introduction
An important issue of dissipative quantum engineering is the NESS dimension analysis. In the case of the GKSL equation, a series of theorems were proved defining algebraic conditions for the Hamiltonian of the system and jump operators for which the stationary solution is unique [41, 42, 43, 44, 45, 46, 47, 48]. Further interest in the formation of dark spaces led to the emergence of theorems that set sufficient conditions for existence of degenerate NESS. So, for example, it was shown that if the GKSL equation has the so-called strong symmetries, given the existence of unitary operators with which both the Hamiltonian of the system and the jump operators commute, the number of NESS is determined by the number of invariant subspaces of operators of such symmetries [49, 50, 51, 52, 48]. An analog of the Lieb-Schultz-Mattis theorem has recently been proved for open systems with translational invariance and strong $U(1)$ symmetry, which states that it is impossible to implement a single NESS with an odd total charge of the system [53]. We emphasize that the described necessary conditions for the existence of degenerate NESS require a special class of dissipative fields: either reflecting the symmetry of the Hamiltonian, or corresponding to a homogeneous system preserving the total charge.
@@
The outline of the paper is as follows. We formulate our dissipative model in Sec. 2. In Sec. 3 we review the third quantization method in application to the considered model. The conditions for existence of ZKM as well as corresponding strong symmetry are derived in Sec. 4. Our general results applied to Kitaev chain are described in Sec. 5. In Sec. 6 we end the paper with conclusions. Some technical details are delegated to the Appendices.
@@ 2.1 The model of dissipative dynamics
In this section, we formulate a model of superconductors that are affected by dissipative fields that do not preserve the number of fermions. As we demonstrate in Appendix 7.1, coupling to reservoirs can be also described as the action of the dissipative fields. Therefore, in what follows, we do not consider coupling to the reservoirs separately. We assume that the evolution of the density matrix of the system affected by $N_{B}$ dissipative channels satisfies the following GKSL equation:
@@
$\hat{w}_{2l-1}=\hat{c}_{l}+\hat{c}^{+}_{l},\qquad\hat{w}_{2l}={\color[rgb]{0,0,0}i(\hat{c}^{+}_{l}-\hat{c}_{l})},\qquad l=1,\dots,N.$ (3)
@@ 2.2 Solution for the unitary evolution
$\mathcal{W}=[\,\underline{\chi}_{1},\penalty 10000\ \underline{\chi}_{2},\ldots,\penalty 10000\ \underline{\chi}_{2a-1},\penalty 10000\ \underline{\chi}_{2a},\ldots,\penalty 10000\ \underline{\chi}_{2N-1},\penalty 10000\ \underline{\chi}_{2N}\,]\,\in\mathbb{R}^{2N\times 2N}.$ (7)
@@
$\varepsilon_{a}={\color[rgb]{0,0,0}\underline{\chi}_{2a-1}^{T}A\,\underline{\chi}_{2a}},\qquad a=1,\dots,N$ (8)
@@ 3.1 The Fock-Liouville space
$|\,\hat{P}_{\underline{\alpha}}\,\rangle=\hat{P}_{\alpha_{1},\alpha_{2},\ldots,\alpha_{2N}}=2^{-N}\,\hat{w}_{1}^{\alpha_{1}}\,\hat{w}_{2}^{\alpha_{2}}\dots\hat{w}_{2N}^{\alpha_{2N}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha_{j}\in\{0,\,1\},$ (13)
@@
[…]
@@ 3.2 Mapping of GKSL equation to the Fock-Liouville space
$\mathcal{U}={\color[rgb]{0,0,0}\begin{pmatrix}I_{N}&I_{N}\\ -iI_{N}&iI_{N}\\ \end{pmatrix}}$. (23)
@@ 3.3 Spectrum of the Liouvillian
$XF+FX^{T}-Y=0\,,\penalty 10000\ \penalty 10000\ Y=-Y^{*}=-Y^{T}.$ (25)
[…]
@@
$\!\!\!\!\!\!{V}\,{\mathcal{L}}\,{V}^{-1}=\begin{pmatrix}B&0\\ 0&-B\end{pmatrix}\,,\quad{V}=\begin{pmatrix}U^{-1}&U^{-1}F\\ 0&U^{T}\\ \end{pmatrix}\,,\penalty 10000\ {V}^{-1}=\begin{pmatrix}U&-F\,(\,U^{-1}\,)^{T}\\ \!0&\,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ (\,U^{-1}\,)^{T}\\ \end{pmatrix}$. (26)
@@
$\check{\mathcal{L}}=-{4}\sum_{l=1}^{2N}\,\beta_{l}\,\check{b}_{l}^{\times}\check{b}_{l}\,;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \begin{pmatrix}\underline{\check{b}}\\ \,\underline{\check{{b}}}^{\times}\\ \end{pmatrix}={V}\begin{pmatrix}\underline{\check{c}}\\ \underline{\check{{c}}}^{+}\end{pmatrix}=\begin{pmatrix}U^{-1}\,\underline{\check{c}}\,+\,U^{-1}\,F\,\underline{\check{c}}^{+}\\ \,U^{T}\,\underline{\check{{c}}}^{+}\end{pmatrix}\,$. (27)
Here the third quantized operators $\check{b}$ and $\check{b}^{\times}$ satisfy fermionic commutation relations: $\{\,\check{b}_{l},\,\check{b}_{m}\,\}=\{\,\check{b}^{\times}_{l},\,\check{b}^{\times}_{m}\,\}=0,\penalty 10000\ \penalty 10000\ \{\,\check{b}_{l},\,\check{b}^{\times}_{m}\,\}=\delta_{lm};\penalty 10000\ \penalty 10000\ l,m=1,\ldots,2N.$ We note that $\check{b}^{\times}_{l}\neq\check{b}_{l}^{+}$.
@@
[…]
@@
[…]
@@ 3.4 Types and robustness of zero modes in closed and open systems
[…]
@@
[…]
@@
[…]
@@
Thus, we argued that occupation of a single ZKM with $\beta_{r}=0$ leads to an indistinguishable change in the structure of a degenerate non-equilibrium stationary state from the point of view of single-particle physical observables. This property of ZKM is different from the property of zero energy states in the isolated case described in Sec. 2. For the latter it holds only for local operators. Such a difference is probably explained by the fact that the NESS forms as a result of hybridization with the environment. In systems with the multiple ZKM, the hybridization behavior becomes more subtle. Nevertheless, the NESS states remain indistinguishable to at least local observables. Therefore, a knowledge of the number of ZKM is crucial for fixing the dimensionality of the dark space in the presence of dissipation.
@@ 3.5 Limit of vanishing dissipation
where we used Eqs. (10) and (24), and the relations $\hat{\rho}^{(1)}\to 0$, $\penalty 10000\ \underline{\psi}_{2l-1(2l)}\to\underline{\chi}_{2l-1(2l)}$, $\sum_{l}\,U_{l\bar{p}}\,\hat{w}_{l}/2\to\hat{\alpha}_{p}^{+}$ in the limit $\gamma{\to}0^{+}$. Enforcing consistency between Eqs. (39)-(41) yields  [^4] Note that in the limit $\gamma\to 0^{+}$, the matrices $F$ and $X$ have a common eigenbasis defined by the matrix $U$. This corresponds to the situation where the so-called spectral and purity gaps coincide [70]. For finite $\gamma$, however, this property generally no longer holds, and the two gaps characterize distinct features of the system.:
@@ 4 The condition for the existence of zero kinetic modes
4.1 Main result: the number and structure of ZKM
Having discussed certain features of zero kinetic modes in the observables of a dissipative system, as well as their relation to Majorana modes in a closed system, we now proceed to formulate the main result concerning the conditions for the formation and the number of ZKMs. We state this result as the following theorem.
Theorem 1.
Consider an isolated system of free fermions with a particle-hole symmetric excitation spectrum with energies $0\leq\varepsilon_{1}\leq\ldots\leq\varepsilon_{N}$. Under the introduction of dissipation via linear Lindblad fields $\underline{l}_{v}$ (Eq. (4)–(5)), the system exhibits zero kinetic modes (ZKMs) with energies $\beta=0$. The number $N_{0}$ of such ZKMs is given by: